OpenStudy (kanwal32):

urgent integration of of (cosx-cos2x)/(1-cosx)

OpenStudy (kanwal32):

@Jaynator495 @radar @Michele_Laino

OpenStudy (kanwal32):

@mathmale @Jaynator495 @radar @hartnn @Michele_Laino

OpenStudy (kanwal32):

@zepdrix

OpenStudy (michele_laino):

hint: we can simplify the function, like below: \[\Large \frac{{\left( {2\cos x + 1} \right)\left( {\cos x - 1} \right)}}{{\left( {\cos x - 1} \right)}} = 2\cos x + 1\]

OpenStudy (kanwal32):

ok thanx

OpenStudy (michele_laino):

so it is suffice to solve this: \[\huge \int {\left( {2\cos x + 1} \right)} \;dx\]